# Products in spinc-cobordism

### v1

@@ 1.1. Spinc cobordism
The spinc-cobordism ring, written $\Omega^{Spin^{c}}_{*}$, is the ring of spinc-cobordism classes of compact smooth spinc-manifolds. The addition is given by disjoint union of manifolds, while the multiplication is Cartesian product. There are several reasons to care about spinc-cobordism: aside from its applications to mathematical physics, e.g. [7] and [28], spinc-cobordism is of particular interest because it is one of the complex-oriented cobordism theories, and consequently there exists a one-dimensional group law on $\Omega^{Spin^{c}}_{*}$ which describes how the first Chern class in spinc-cobordism behaves on a tensor product of complex line bundles. See [1] or [12] for these classical ideas, whose consequences for complex cobordism (as in [20]) have been enormous, but whose consequences for spinc-cobordism have apparently never been fully explored [^1] In future work, the authors hope to apply the results about the ring structure of the spinc-cobordism ring obtained in this paper to the problem of describing the formal group law on the spinc-cobordism ring in formal-group-law theoretic terms, similar to what Quillen did for complex and unoriented cobordism in [16],[17], what Baker–Morava did for $2$-inverted symplectic bordism in [6], and what Buchstaber did for symplectic bordism [9]. It seems impossible to get much understanding of the formal group law of spinc-cobordism without first coming to some understanding of the structure of its coefficient ring $\Omega^{Spin^{c}}_{*}$, which is the goal of this paper..
Since spinc-cobordism is an example of a “$(B,f)$-cobordism theory” in the sense of Thom, the general results of [25] ensure that there exists a spectrum $MSpin^{c}$ such that $\pi_{*}(MSpin^{c})\cong\Omega^{Spin^{c}}_{*}$. The homotopy type of $MSpin^{c}$ is understood as follows.
@@
The map $\pi:BSpin^{c}\longrightarrow BSO\times K(\mathbb{Z},2)$ is an odd-primary homotopy equivalence and induces an isomorphism $\Omega^{Spin^{c}}_{*}[\frac{1}{2}]\cong\Omega^{SO}_{*}(K(\mathbb{Z},2))[\frac{1}{2}]$, and consequently $MSpin^{c}[\frac{1}{2}]\cong MSO[\frac{1}{2}]\wedge\mathbb{C}P^{\infty}$. Since $MSpin^{c}$ is complex-oriented, we have an isomorphism of graded rings $\pi_{*}(MSpin^{c}[\frac{1}{2}])\cong\pi_{*}(MSO)[\frac{1}{2}][[x]]$ with $x\in\pi_{2}(MSpin^{c})$. The ring $\pi_{*}(MSO)[\frac{1}{2}]\cong\pi_{*}(MSp)[\frac{1}{2}]\cong\pi_{*}(MSpin)[\frac{1}{2}]$ is understandable in various ways; see Wall [26] or Baker–Morava [6] for different approaches.
@@ 1.2. The mod $2$ spinc-cobordism ring in low degrees
A traditional notation for the unoriented bordism ring $\Omega^{O}_{*}\cong MO_{*}$ is $\mathfrak{N}_{*}$. In the 1968 book [24, pg. 351], Stong asks:
Open question: Can one determine these images nicely as subrings of $\mathfrak{N}_{*}$?
By “these images,” Stong refers to the images of the natural maps $\Omega^{Spin}_{*}\rightarrow\mathfrak{N}_{*}$ and $\Omega^{Spin^{c}}_{*}\rightarrow\mathfrak{N}_{*}$. Our approach to understanding the mod $2$ spinc-cobordism ring begins by answering Stong’s open question in a range of degrees. We use the Anderson–Brown–Peterson splitting [2], product structure in the Adams spectral sequences, and Thom’s determination of $\mathfrak{N}_{*}$ using symmetric polynomials [25] to develop a method for calculating the image of the map $\Omega^{Spin^{c}}_{*}\rightarrow\mathfrak{N}_{*}$ through degree $d$, for any fixed choice of integer $d$. Our method gives a presentation for $\Omega^{Spin^{c}}/(2,\beta)$ through degree $d$, since the map $\Omega^{Spin^{c}}/(2,\beta)\longrightarrow\mathfrak{N}_{*}$ is injective. We carry out computer calculation using our method to obtain our first main theorem:
@@
Consider the spinc-cobordism ring as a graded algebra over the graded ring $S:=\mathbb{Z}_{(2)}[\beta,Z_{2j}:j+1\mbox{ not\ a\ power\ of\ }2]/(\beta Z_{2j},2Z_{2j}\mbox{\ for\ odd\ }j).$ Let $J$ be the ideal of $S$ generated by all the elements $Z_{2j}$ with $j$ odd. Then $J$ embeds, as a non-unital graded $S$-algebra, into the $2$-torsion ideal $\pi_{*}(Z)$ of the spinc-cobordism ring.
@@
We also calculate the image of the map $\Omega^{Spin}_{*}\rightarrow\mathfrak{N}_{*}$ through degree 31 in Proposition 3.2 via a similar method to the one used to calculate the image of the map $\Omega^{Spin^{c}}_{*}\rightarrow\mathfrak{N}_{*}$ in Theorem A. There is a noteworthy geometric consequence of Proposition 3.2. In the 1965 paper [14], Milnor asks this question:
@@
Here $s_{6}$ is a certain symmetric polynomial, and $p_{1},\dots,p_{6}$ are Pontryagin classes. The reason for Milnor’s question is that, in [14], Milnor proves that, for a compact smooth manifold $M$ of dimension $\leq 23$, the following conditions are equivalent:
@@
Milnor points out that, if there exists a compact spin manifold $\Sigma$ whose Pontryagin number $s_{6}(p_{1},\dots,p_{6})[\Sigma]$ is odd, then these conditions would fail to be equivalent in dimension $24$. Anderson–Brown–Peterson [2],[3] established that, as a consequence of their splitting of $2$-local $MSpin$, there does indeed exist such a compact spin manifold $\Sigma$. However, it seems that no explicit description of that $24$-dimensional compact spin manifold has been given in the literature (or anywhere else, as far as we know).
In Theorem 3.6, we give an explicit formula for the unoriented bordism class of such a compact spin manifold $\Sigma$, as a disjoint union of products of real projective spaces and squares of Dold manifolds. We refer the reader to the Theorem 3.6 for a statement of that formula, which is lengthy. The formula is obtained using our calculation of the image of the map $\Omega^{Spin}_{24}\rightarrow\mathfrak{N}_{24}$ and the manifold representatives calculated in Proposition 3.5.
@@ 1.4. Determination of the mod $2$ spinc-cobordism ring up to inseparable isogeny
Thom’s famous calculation [25] established that the unoriented bordism ring $\Omega^{O}_{*}=\mathfrak{N}_{*}$ is isomorphic to a polynomial algebra over $\mathbb{F}_{2}$. A theorem of Stong [23, Proposition 14] shows that the spinc cobordism ring, reduced modulo torsion and then reduced modulo $2$, is also isomorphic to a polynomial $\mathbb{F}_{2}$-algebra.
@@
with $\beta$ the Bott element in degree $2$, with $y_{4i}$ in degree $4i$, and with $Z_{4j-2}$ in degree $24j-$ 2.
@@
  • and a generator in degree $24j-$ 2 for all positive integers $j$ such that $j$ is not a power of $2$.
@@ 1.6. Acknowledgements
The first author would like to thank Bob Bruner for many helpful conversations related to this work, and the Simons Foundation for providing the license for a copy of Magma [8] used in calculations. The first author was partially supported by the electronic Computational Homotopy Theory (eCHT) research community, funded by National Science Foundation Research Training Group in the Mathematical Sciences grant 2135884.
@@ 2. Preliminaries
2.1. Review of the cohomology of the spectra $MSpin^{c}$ and $MSpin$
@@
where $I$ is the ideal $\langle w_{3},\operatorname{{\rm Sq}}^{2}(w_{3}),\operatorname{{\rm Sq}}^{2}(Sq^{4}(w_{3})),\operatorname{{\rm Sq}}^{8}(Sq^{4}(Sq^{2}(w_{3})),...\rangle$.
@@
where $J$ is the ideal $\langle w_{2},w_{3},\operatorname{{\rm Sq}}^{2}(w_{3}),\operatorname{{\rm Sq}}^{2}(Sq^{4}(w_{3})),\operatorname{{\rm Sq}}^{8}(Sq^{4}(Sq^{2}(w_{3})),...\rangle$.
@@
    which has the fewest nonzero monomial terms among all those which have $X^{a_{1}}X^{a_{2}}\dots X^{a_{n}}$ as a monomial term.
@@
See [24], particularly pages 71 and 96 and surrounding material, for a nice exposition of the following result, which dates back to Thom [25]: let $\Lambda$ be the set of unordered finite-length tuples of positive integers, each of which is not equal to $12^{a}-$ 1 for any integer $a$. Such integers are called “non-dyadic,” and such partitions are called “non-dyadic partitions.” For each $\lambda\in\Lambda$, write $\left|\lambda\right|$ for the length of $\lambda$, and write $\left|\left|\lambda\right|\right|$ for the sum of the elements of $\lambda$. Consider the polynomial
@@
(5) $\left\{P_{\lambda}(w_{1},\dots,w_{\left|\lambda\right|})U:\lambda\in\Lambda,\left|\left|\lambda\right|\right|=n\right\}.$
@@
Consequently, if we write $\{Y_{\lambda}\}$ for the basis of $\pi_{*}(MO)$ dual to the basis (5), then $Y_{\lambda}Y_{\lambda^{\prime}}=Y_{\lambda\coprod\lambda^{\prime}}$, and $\mathfrak{N}_{*}\cong\pi_{*}(MO)\cong\mathbb{F}_{2}[Y_{(2)},Y_{(4)},Y_{(5)},Y_{(6)},Y_{(8)},\dots]$, with $Y_{(i)}$ in degree $i$. We will sometimes write $Y_{i}$ as an abbreviation for Thom’s generator $Y_{(i)}$ of $\mathfrak{N}_{*}$.
@@ 2.3. Maps between bordism theories
$BString=BO\langle 8\rangle\longrightarrow BSpin=BO\langle 4\rangle\longrightarrow BSO=BO\langle 2\rangle\longrightarrow BO.$
@@
${K(\mathbb{Z}/2\mathbb{Z},0)}$${Spin(n)}$${SO(n)}$${K(\mathbb{Z},1)}$${Spin^{c}(n)}$${SO(n)}$${K(\mathbb{Z},1)\cong U(1)}$${U(1)}${*}
@@
(6) $MString\longrightarrow MSpin\longrightarrow MSpin^{c}\longrightarrow MSO\longrightarrow MO.$
@@
Hence our broad strategy for calculating $\Omega^{Spin^{c}}_{*}/(2,\beta)\cong MSpin^{c}_{*}/(2,\beta)$ and $\Omega^{Spin}_{*}/(2,\eta,\alpha,\beta)\cong MSpin_{*}/(2,\eta,\alpha,\beta)$, and the natural maps $MSpin_{*}/(2,\eta,\alpha,\beta)\rightarrow MSpin^{c}_{*}/(2,\beta)\rightarrow MO_{*}$, is to calculate the $A_{*}$-comodule primitives in $H_{*}(MSpin^{c};\mathbb{F}_{2})$ and in $H_{*}(MSpin;\mathbb{F}_{2})$, regarding each as $A_{*}$-subcomodule algebras of $H_{*}(MO;\mathbb{F}_{2})\cong\mathbb{F}_{2}[w_{1},w_{2},w_{3},\dots]\{U\}$. The resulting information will describe $\Omega^{Spin^{c}}_{*}/(2,\beta)$ as a subring of $\mathfrak{N}_{*}$. Details of this strategy are given in the description of the computational method in the proof of Proposition 3.1.
@@
[…]
@@ 3. The $Spin^{c}$ bordism ring in low degrees
In the statement of Proposition 3.1, we use Thom’s presentation $\mathbb{F}_{2}[Y_{2},Y_{4},Y_{5},...]$ for the unoriented cobordism ring $\mathfrak{N}_{*}$.
@@
The image of the map $\Omega^{Spin^{c}}_{*}\longrightarrow\mathfrak{N}_{*}$ agrees, in degrees $\leq 33$, with subring of $\mathfrak{N}_{*}$ generated by the elements
@@
Proof.
@@
The induction begins at degree 00, where there is nothing to say: the empty partition $\varnothing$ yields the unique $A$-module primitive $\varnothing\cdot U$ in $H^{0}(MSpin^{c};\mathbb{F}_{2})$. For each integer $n\in[1,d]$, the product $\operatorname{{\rm Sq}}^{n}\cdot\varnothing\cdot U\in H^{n}(MSpin^{c};\mathbb{F}_{2})$ is simply $w_{n}U$ modulo the relations in $H^{*}(MSpin^{c};\mathbb{F}_{2})$, by the classical formula $\operatorname{{\rm Sq}}^{n}U=w_{n}U$ for the action of Steenrod squares on the Thom class in $H^{0}(MO;\mathbb{F}_{2})$. Record the elements $\{\operatorname{{\rm Sq}}^{1}U,\operatorname{{\rm Sq}}^{2}U,\dots,\operatorname{{\rm Sq}}^{d}U\}$ in an unordered list $D$. Here the symbol $D$ stands for “decomposable,” as we will use it to build up a list of $A$-module decomposables in $H^{*}(MSpin^{c};\mathbb{F}_{2})$ in degrees $\leq d$.
We are not done with the initial step in the induction: for each nonzero element $\operatorname{{\rm Sq}}^{n}U$, we calculate $\{\operatorname{{\rm Sq}}^{1}\operatorname{{\rm Sq}}^{n}U,\operatorname{{\rm Sq}}^{2}\operatorname{{\rm Sq}}^{n}U,\dots,\operatorname{{\rm Sq}}^{d-n}\operatorname{{\rm Sq}}^{n}U\}$ using the Thom formula $\operatorname{{\rm Sq}}^{n}U=w_{n}U$ and the Wu formula ([27], but see [15, pg. 94] for a textbook reference) for the action of $\operatorname{{\rm Sq}}^{n}$ on Stiefel–Whitney classes, and we include the results in $D$. Now $D$ contains an $\mathbb{F}_{2}$-linear basis for the $A$-submodule of $H^{*}(MSpin^{c};\mathbb{F}_{2})$ generated by the Thom class $U$, in all degrees $\leq d$.
@@
We have produced a list $G$ of $\mathbb{F}_{2}$-linear combinations of non-dyadic partitions of degree $<n$, such that the $\mathbb{F}_{2}$-linear span of $\{P_{\lambda}U:\lambda\in G\}$ is a basis for the set of $A$-module primitives in $H^{*}(MSpin^{c};\mathbb{F}_{2})$ in all degrees $<n$. We have also produced a list $D$ of $\mathbb{F}_{2}$-linear combinations of non-dyadic partitions of degree $\leq d$, such that the $\mathbb{F}_{2}$-linear span of $\{P_{\lambda}U:\lambda\in D\}$ is precisely the $A$-submodule of $H^{*}(MSpin^{c};\mathbb{F}_{2})$ generated by $G$ in degrees $\leq d$.
@@
Once we complete the $n=d$ step, we have an $\mathbb{F}_{2}$-linear basis for the image of the map $\Omega^{Spin^{c}}_{*}\rightarrow\mathfrak{N}_{*}$ in all degrees $\leq d$, expressed in terms of Thom’s partition basis for $\mathfrak{N}_{*}$. Consequently we have a description of $\Omega^{Spin^{c}}_{*}/(2,\beta)$, in all degrees $\leq d$, as a subring of $\mathfrak{N}_{*}\cong\mathbb{F}_{2}[Y_{2},Y_{4},Y_{5},Y_{6},Y_{8},Y_{9},\dots]$. ∎
In principle there is no obstruction to using the same method to make calculations of products in $\Omega^{Spin^{c}}_{*}$ in degrees $>33$. We stopped at degree $33$ simply because, around the time we completed degree $33$, we could see enough of the ring structure of $\Omega^{Spin^{c}}_{*}$ to prove that the mod $(2,\beta)$ spinc-cobordism ring is not a polynomial algebra, and to suggest the right statements for Proposition 4.2 and Theorem C. The products in $\Omega^{Spin^{c}}_{*}$ required for the proof of Theorem 3.6 are known as soon as one computes the ring structure through degree $24$.

### v2

@@ 1.1. Spinc cobordism
The spinc-cobordism ring, written $\Omega^{Spin^{c}}_{*}$, is the ring of spinc-cobordism classes of compact smooth spinc-manifolds. The addition is given by disjoint union of manifolds, while the multiplication is Cartesian product. There are several reasons to care about spinc-cobordism: aside from its applications to mathematical physics, e.g. [7] and [29], spinc-cobordism is of particular interest because it is one of the complex-oriented cobordism theories, and consequently there exists a one-dimensional group law on $\Omega^{Spin^{c}}_{*}$ which describes how the first Chern class in spinc-cobordism behaves on a tensor product of complex line bundles. See [1] or [13] for these classical ideas, whose consequences for complex cobordism (as in [21]) have been enormous, but whose consequences for spinc-cobordism have apparently never been fully explored.
Since spinc-cobordism is an example of a “$(B,f)$-cobordism theory” in the sense of Thom, the general results of [26] ensure that there exists a spectrum $MSpin^{c}$ such that $\pi_{*}(MSpin^{c})\cong\Omega^{Spin^{c}}_{*}$. The homotopy type of the spectrum $MSpin^{c}$ is understood as follows.
@@
The map $\pi:BSpin^{c}\longrightarrow BSO\times K(\mathbb{Z},2)$ is an odd-primary homotopy equivalence and induces an isomorphism $\Omega^{Spin^{c}}_{*}[\frac{1}{2}]\cong\Omega^{SO}_{*}(K(\mathbb{Z},2))[\frac{1}{2}]$, and consequently $MSpin^{c}[\frac{1}{2}]\cong MSO[\frac{1}{2}]\wedge\mathbb{C}P^{\infty}$.
@@
Remark 1.1.
In principle, the ring structure on $\Omega^{Spin^{c}}_{*}$ away from $2$ is understood, although only in a rather indirect way. Here is how it works: from the complex-orientability of $MSO$, one gets an isomorphism of rings $MSO[\frac{1}{2}]^{*}(\mathbb{C}P^{\infty})\cong MSO[\frac{1}{2}]^{*}[[X]]$. The ring $MSO[\frac{1}{2}]^{*}[[X]]$ is also the “covariant bialgebra” of the formal group law of $MSO[\frac{1}{2}]_{*}$, in the sense of [12, Chapter 36]. Hence one can use the formal group law on $MSO[\frac{1}{2}]^{*}$ (whose universal property is given by [6]) to understand the coproduct on $MSO[\frac{1}{2}]^{*}[[X]]$, whose dual, in an appropriate sense, is responsible for the ring structure on $MSO[\frac{1}{2}]_{*}(\mathbb{C}P^{\infty})\cong\Omega^{Spin^{c}}_{*}[\frac{1}{2}]$. This gives a means of understanding the ring $\Omega^{Spin^{c}}_{*}[\frac{1}{2}]$, although we know of nowhere in the literature where this has been carried out in any further detail.
@@ 1.2. The mod $2$ spinc-cobordism ring in low degrees
There are natural forgetful maps, $\Omega^{Spin}_{*}\rightarrow\Omega^{O}_{*}$ and $\Omega^{Spin^{c}}_{*}\rightarrow\Omega^{O}_{*}$, from the spin and spinc-cobordism rings to the unoriented cobordism ring. Since $\Omega^{O}_{*}$ is well-understood, it is natural to try to understand the images of these forgetful maps. In the 1968 book [25, pg. 351], Stong asks:
Open question: Can one determine these images nicely as subrings of $\Omega^{O}_{*}$?
Our approach to understanding the mod $2$ spinc-cobordism ring begins by answering Stong’s open question in a range of degrees. We use the Anderson–Brown–Peterson splitting [2], product structure in the Adams spectral sequences, and Thom’s determination of $\Omega^{O}_{*}$ using symmetric polynomials [26] to develop a method for calculating the image of the map $\Omega^{Spin^{c}}_{*}\rightarrow\Omega^{O}_{*}$ through degree $d$, for any fixed choice of integer $d$. Our method gives a presentation for $\Omega^{Spin^{c}}/(2,\beta)$ through degree $d$, since the map $\Omega^{Spin^{c}}/(2,\beta)\longrightarrow\Omega^{O}_{*}$ is injective. We carry out computer calculation using our method to obtain our first main theorem:
@@
[…]
@@
We have emphasized that the structure of the spinc-cobordism ring has remained mysterious, despite the great control over its additive structure which comes from the Anderson–Brown–Peterson splitting. In these respects, spinc-bordism is not unique: the structure of the spin-bordism ring is also quite mysterious, despite Anderson–Brown–Peterson [2],[3] also constructing a splitting of the $2$-local spin bordism spectrum $MSpin$ into copies of $H\mathbb{F}_{2}$, $ko_{(2)}$, and a suitably-connected cover of $ko_{(2)}$. Our methods were developed for spinc-cobordism, but in this paper we also apply them to spin cobordism. In Proposition 3.2, we obtain a calculation of the image of the map $\Omega^{Spin}_{*}\rightarrow\Omega^{O}_{*}$ through degree 31. This has a noteworthy geometric consequence, which we will now describe.
@@
In the 1965 paper [15], Milnor asks this question:
@@
Here $s_{6}$ is a certain symmetric polynomial (the sixth Girard–Newton polynomial, which converts from the elementary symmetric polynomials to the power sum polynomials), and $p_{1},\dots,p_{6}$ are Pontryagin classes. The reason for Milnor’s question is that, in [15], Milnor proves that, for a compact smooth manifold $M$ of dimension $\leq 23$, the following conditions are equivalent:
@@
Milnor points out that, if there exists a compact spin manifold $\Sigma$ whose Pontryagin number $s_{6}(p_{1},\dots,p_{6})[\Sigma]$ is odd, then these conditions would fail to be equivalent in dimension $24$. Anderson–Brown–Peterson [2],[3] established that, as a consequence of their splitting of $2$-local MSpinMSpin, there does indeed exist such a compact spin manifold $\Sigma$. However, it seems that no explicit description of that $24$-dimensional compact spin manifold has been given in the literature (or anywhere else, as far as we know).
In Theorem 3.7, we give an explicit formula for the unoriented bordism class of such a compact spin manifold $\Sigma$, as a disjoint union of products of real projective spaces and squares of Dold manifolds. We refer the reader to the Theorem 3.7 for a statement of that formula, which is lengthy. The formula is obtained using our calculation of the image of the map $\Omega^{Spin}_{24}\rightarrow\Omega^{O}_{24}$ and the manifold representatives calculated in Proposition 3.6.
@@ 1.4. Determination of the mod $2$ spinc-cobordism ring up to inseparable isogeny
Thom’s famous calculation [26] established that the unoriented bordism ring $\Omega^{O}_{*}$ is isomorphic to a polynomial algebra over $\mathbb{F}_{2}$. A theorem of Stong [24, Proposition 14] shows that the spinc cobordism ring, reduced modulo torsion and then reduced modulo $2$, is also isomorphic to a polynomial $\mathbb{F}_{2}$-algebra.
@@
with $\beta$ the Bott element in degree $2$, with $y_{4i}$ in degree $4i$, and with $Z_{4j-2}$ in degree 4j24j-2.
@@
  • and a generator in degree 4j24j-2 for all positive integers $j$ such that $j$ is not a power of $2$.
@@ 1.5. Conventions
1.6. Funding
The first author was partially supported by the electronic Computational Homotopy Theory (eCHT) research community, funded by National Science Foundation Research Training Group in the Mathematical Sciences grant 2135884.
@@ 1.7. Acknowledgements
The first author would like to thank Bob Bruner for many helpful conversations related to this work, and the Simons Foundation for providing the license for a copy of Magma [8] used in calculations. We are also grateful to an anonymous referee for helpful comments.
@@ 2. Preliminaries
2.1. Review of the cohomology of the spectra $MSpin^{c}$ and MSpinMSpin
@@
where $I$ is the ideal $\langle w_{3},\operatorname{{\rm Sq}}^{2}(w_{3}),\operatorname{{\rm Sq}}^{2}(Sq^{4}(w_{3})),\operatorname{{\rm Sq}}^{8}(Sq^{4}(Sq^{2}(w_{3})),$$\rangle.$
@@
where $J$ is the ideal $\langle w_{2},w_{3},\operatorname{{\rm Sq}}^{2}(w_{3}),\operatorname{{\rm Sq}}^{2}(Sq^{4}(w_{3})),\operatorname{{\rm Sq}}^{8}(Sq^{4}(Sq^{2}(w_{3})),$$\rangle.$
@@
    which has the fewest nonzero monomial terms among all those which have $X_{1}^{a_{1}}X_{2}^{a_{2}}\dots X_{n}^{a_{n}}$ as a monomial term.
@@
[…]
@@
[…]
@@
Consequently, if we write $\{Y_{\lambda}\}$ for the basis of $\pi_{*}(MO)$ dual to the basis (5), then $Y_{\lambda}Y_{\lambda^{\prime}}=Y_{\lambda\coprod\lambda^{\prime}}$, and $\Omega^{O}_{*}\cong\pi_{*}(MO)\cong\mathbb{F}_{2}[Y_{(2)},Y_{(4)},Y_{(5)},Y_{(6)},Y_{(8)},\dots]$, with $Y_{(i)}$ in degree $i$. We will sometimes write $Y_{i}$ as an abbreviation for Thom’s generator $Y_{(i)}$ of $\Omega^{O}_{*}$.
@@ 2.3. Maps between bordism theories
$BSpin=BO\langle 4\rangle\longrightarrow BSO=BO\langle 2\rangle\longrightarrow BO.$
@@
[…]
@@
(6) $MSpin\longrightarrow MSpin^{c}\longrightarrow MSO\longrightarrow MO.$
@@
Hence our broad strategy for calculating $\Omega^{Spin^{c}}_{*}/(2,\beta)\cong MSpin^{c}_{*}/(2,\beta)$ and $\Omega^{Spin}_{*}/(2,\eta,\alpha,\beta)\cong MSpin_{*}/(2,\eta,\alpha,\beta)$, and the natural maps $MSpin_{*}/(2,\eta,\alpha,\beta)\rightarrow MSpin^{c}_{*}/(2,\beta)\rightarrow MO_{*}$, is to calculate the $A_{*}$-comodule primitives in $H_{*}(MSpin^{c};\mathbb{F}_{2})$ and in $H_{*}(MSpin;\mathbb{F}_{2})$, regarding each as $A_{*}$-subcomodule algebras of $H_{*}(MO;\mathbb{F}_{2})$. The resulting information will describe $\Omega^{Spin^{c}}_{*}/(2,\beta)$ as a subring of $\Omega^{O}_{*}$. Details of this strategy are given in the description of the computational method in the proof of Proposition 3.1.
@@
${\text{MSpin}}_{*}/(2,\eta,\alpha,B)$$H_{*}(\text{MSpin};\mathbb{F}_{2})$$\text{MSpin}^{c}_{*}/(2,\beta)$$H_{*}(\text{MSpin}^{c};\mathbb{F}_{2})$$\text{MSO}_{*}/(2)$$H_{*}(\text{MSO};\mathbb{F}_{2})$$\text{MO}_{*}$$H_{*}(\text{MO};\mathbb{F}_{2})$}}
@@ 3. The $Spin^{c}$ bordism ring in low degrees
In the statement of Proposition 3.1, we use Thom’s presentation $\mathbb{F}_{2}[Y_{2},Y_{4},Y_{5},$...] for the unoriented cobordism ring $\Omega^{O}_{*}$.
@@
The image of the map $\Omega^{Spin^{c}}_{*}\longrightarrow\Omega^{O}_{*}$ agrees, in degrees $\leq 33$, with the subring of $\Omega^{O}_{*}$ generated by the elements
@@
Each of the elements $T_{i}$ defined in Proposition 3.1 is a linear combination of monomials in $\Omega^{O}_{*}$. Those monomials are generally not individually members of $\Omega^{Spin^{c}}_{*}$: for example, $Y_{14}Y_{5}^{2}\in\Omega^{O}_{24}$ does not lift to an element of $\Omega^{Spin^{c}}_{24}$, even though a linear combination of $Y_{14}Y_{5}^{2}$ with other monomials in degree $24$ does lift to the element $T_{24}\in\Omega^{Spin^{c}}_{24}$.
Proof of Proposition 3.2.
@@
The induction begins at degree 0, where there is nothing to say: the empty partition $\varnothing$ yields the unique $A$-module primitive $\varnothing\cdot U$ in $H^{0}(MSpin^{c};\mathbb{F}_{2})$. For each integer $n\in[1,d]$, the product $\operatorname{{\rm Sq}}^{n}\cdot\varnothing\cdot U\in H^{n}(MSpin^{c};\mathbb{F}_{2})$ is simply $w_{n}U$ modulo the relations in $H^{*}(MSpin^{c};\mathbb{F}_{2})$, by the classical formula $\operatorname{{\rm Sq}}^{n}U=w_{n}U$ for the action of Steenrod squares on the Thom class in $H^{0}(MO;\mathbb{F}_{2})$. Record the elements $\{\operatorname{{\rm Sq}}^{1}U,\operatorname{{\rm Sq}}^{2}U,\dots,\operatorname{{\rm Sq}}^{d}U\}$ in an unordered list $D$. Here the symbol $D$ stands for “decomposable,” as we will use it to build up a list of $A$-module decomposables in $H^{*}(MSpin^{c};\mathbb{F}_{2})$ in degrees $\leq d$.
We are not done with the initial step in the induction: for each nonzero element $\operatorname{{\rm Sq}}^{n}U$, we calculate $\{\operatorname{{\rm Sq}}^{1}\operatorname{{\rm Sq}}^{n}U,\operatorname{{\rm Sq}}^{2}\operatorname{{\rm Sq}}^{n}U,\dots,\operatorname{{\rm Sq}}^{d-n}\operatorname{{\rm Sq}}^{n}U\}$ using the Thom formula $\operatorname{{\rm Sq}}^{n}U=w_{n}U$ and the Wu formula ([28], but see [16, pg. 94] for a textbook reference) for the action of $\operatorname{{\rm Sq}}^{n}$ on Stiefel–Whitney classes, and we include the results in $D$. We keep going: calculate all the three-fold composites of Steenrod squares applied to $U$ landing in degrees $\leq d$, then all the four-fold composites of Steenrod squares applied to $U$ landing in degrees $\leq d$, and so on. We emphasize the phrase “landing in degrees $\leq d$” because it is what ensures that this calculation eventually terminates!
After we are done with that inductive calculation, $D$ now contains an $\mathbb{F}_{2}$-linear basis for the $A$-submodule of $H^{*}(MSpin^{c};\mathbb{F}_{2})$ generated by the Thom class $U$, in all degrees $\leq d$.
@@
We have produced a list $G$ of $\mathbb{F}_{2}$-linear combinations of non-dyadic partitions of degree <n<n, such that the $\mathbb{F}_{2}$-linear span of $\{P_{\lambda}U:\lambda\in G\}$ is a basis for the set of $A$-module primitives in $H^{*}(MSpin^{c};\mathbb{F}_{2})$ in all degrees <n<n. We have also produced a list $D$ of $\mathbb{F}_{2}$-linear combinations of non-dyadic partitions of degree $\leq d$, such that the $\mathbb{F}_{2}$-linear span of $\{P_{\lambda}U:\lambda\in D\}$ is precisely the $A$-submodule of $H^{*}(MSpin^{c};\mathbb{F}_{2})$ generated by $G$ in degrees $\leq d$.
@@
Once we complete the $n=d$ step, we have an $\mathbb{F}_{2}$-linear basis for the image of the map $\Omega^{Spin^{c}}_{*}\rightarrow\Omega^{O}_{*}$ in all degrees $\leq d$, expressed in terms of Thom’s partition basis for $\Omega^{O}_{*}$. Consequently we have a description of $\Omega^{Spin^{c}}_{*}/(2,\beta)$, in all degrees $\leq d$, as a subring of $\Omega^{O}_{*}\cong\mathbb{F}_{2}[Y_{2},Y_{4},Y_{5},Y_{6},Y_{8},Y_{9},\dots]$. ∎
In principle there is no obstruction to using the same method to make calculations of products in $\Omega^{Spin^{c}}_{*}$ in degrees >33>33. We stopped at degree $33$ simply because, around the time we completed degree $33$, we could see enough of the ring structure of $\Omega^{Spin^{c}}_{*}$ to prove that the mod $(2,\beta)$ spinc-cobordism ring is not a polynomial algebra, and to suggest the right statements for Proposition 4.2 and Theorem C. The products in $\Omega^{Spin^{c}}_{*}$ required for the proof of Theorem 3.7 are known as soon as one computes the ring structure through degree $24$. Completing the calculation through degree 33 took more than a week on consumer hardware. We anticipate that a calculation of even a few degrees beyond 33 will potentially take months to complete without the use of high-performance computing hardware or an improved algorithm.